# Mitigating Malicious Attacks in Federated Learning via Confidence-aware Defense

### v1

@@ Introduction
Mild (25%) FedAvg 68.32 68.60 49.84 52.89 55.40 64.37 49.44 59.20
MKrum 66.84 66.25 29.94 28.12 44.59 36.90 57.16 43.17
TrimMean 65.20 65.20 31.63 34.22 53.69 61.35 51.58 60.33
AFA 68.88 69.09 69.09 68.80 68.89 68.71 68.99 69.17
FLDetector 39.74 46.32 61.94 56.92 32.66 32.37 25.93 35.61
DeFL 67.95 68.57 48.02 51.51 65.67 69.04 62.97 68.74
CAD (Ours) 69.36 69.47 69.67 69.18 69.85 69.43 69.94 69.43
Moderate (50%) FedAvg 62.91 62.70 12.94 15.04 19.02 18.37 21.40 22.25
MKrum 10.40 11.99 10.15 10.00 11.00 13.93 17.19 17.15
TrimMean 37.26 36.88 11.55 12.35 16.92 17.27 17.49 17.20
AFA 60.32 60.85 10.42 10.56 16.65 17.89 15.42 17.42
FLDetector 34.84 38.81 14.41 - 23.61 - 20.47 -
DeFL 63.02 62.31 23.17 23.95 21.60 - 48.26 20.76
CAD (Ours) 64.27 64.19 64.44 64.96 64.00 65.07 64.75 65.13
Severe (75%) FedAvg 50.51 53.46 10.56 12.08 16.71 17.72 17.79 17.60
MKrum 11.92 10.80 10.03 10.60 13.34 13.84 15.84 15.67
TrimMean 21.17 20.23 10.01 10.42 14.57 15.03 17.52 16.60
AFA 20.23 22.17 10.06 10.51 11.54 11.62 13.84 10.03
FLDetector 18.33 18.04 11.04 12.08 16.71 17.72 17.79 -
DeFL 52.18 52.72 16.76 - 20.11 - 19.42 20.02
CAD (Ours) 57.68 57.79 56.75 56.54 56.96 57.04 56.88 56.48
Baseline Defenses
We compared CAD with state-of-the-art defenses including: FedAvg (McMahan et al. 2017), Mkrum (Blanchard et al. 2017), TrimMean (Yin et al. 2018), AFA (Muñoz-González, Co, and Lupu 2019), FLDetector (Zhang and et al. 2022), DeFL (Yan et al. 2023). We consider two settings regarding the adversary’s knowledge as prior works: (a) Full, where the adversary knows the gradients of benign clients, and (b) Partial, where the adversary is unaware of the gradient updates shared by benign clients.
Malicious Client Attacks
We evaluated CAD model on both model poisoning attacks and data poisoning attacks:
@@
• Little is Enough (LIE): Adds small amounts of noise to the average of benign gradients. The adversary computes the average $\mu$ and standard deviation $\sigma$ of benign gradients, then applies a coefficient $z$ to obtain the malicious update $\mu+z\sigma$.
• Min-Max (MM): Computes malicious gradients such that their maximum distance from any gradient is upper bounded by the maximum distance between any two benign gradients, placing them close to the benign gradient cluster.
• Min-Sum (MS): Ensures the sum of squared distances of malicious gradients from all benign gradients is upper bounded by the sum of squared distances among benign gradients. Malicious gradients are kept identical to maximize impact.
Implementation Details
We employed a total of $N=50$ clients in our experiments. Each client was trained for 20 epochs using stochastic gradient descent to update its local model. The number of training rounds was set to 200 to ensure convergence for all DNN models on all datasets considered in this paper. The learning rate was initialized at 0.01 and decayed by a constant factor of 0.99 after each communication round. The batch size was set to 16, with a weight decay of $1\times 10^{-3}$. The percentage of malicious clients was set to three levels: mild (10%), moderate (25%), and severe (75%), to illustrate the robustness of our CAD model. For data poisoning attacks, confidence scores were directly obtained from the model’s output probabilities on the training data. For model poisoning attacks, confidence scores were obtained by inferencing on local training data before sending it to the parameter server. We implemented our defenses, in PyTorch on a single NVIDIA RTX A100 GPU.
@@
Comparison on Defense Robustness under Different Attack Intensities
@@
We evaluated the robustness of our CAD model and other state-of-the-art defenses under different levels of attack intensity with the CIFAR-10 dataset and the VGG model. The results, as shown in Table 2, indicate that under mild attacks (25% malicious clients), most defense methods perform reasonably well, with our CAD model achieving the highest accuracy across all attack types. However, as the intensity of the attacks increases to moderate (50%) and severe (75%), the performance of most methods significantly deteriorates. In contrast, our CAD model consistently maintains high accuracy and robustness by outperforming other methods under all attack intensities. This demonstrates the effectiveness of our confidence-based detection and re-weighted aggregation approach in mitigating the impact of various attack types, even under severe attack conditions.
Defense Robustness across Various Models and Datasets
@@
Table 3 compares the performance of CAD with other state-of-the-art defense methods across different models and datasets. The results demonstrate that CAD consistently outperforms other defenses with high accuracy under various attacks. For MNIST with MLP, most defense methods perform well under label shuffling, but CAD achieves the highest accuracy across all attack types. Similarly, for Fashion MNIST with AlexNet, CAD maintains superior performance, especially against LIE and MM attacks, where other methods significantly collapsed in accuracy. In the CIFAR-10 dataset, both with VGG-11 and AlexNet models, the robustness of CAD stands out. While other defenses struggle to converge, particularly under LIE and MM attacks, CAD consistently achieves the best results. This highlights the effectiveness of our method in providing robust defense against various attacks, even under different model and dataset combinations. These results demonstrated that CAD is not only effective in mitigating different types of attacks but also adaptable to various models and datasets, thereby making it a reliable defense mechanism in FL systems.
@@
Table 3: Defense robustness across various models and datasets. Accuracy of different defense methods under various attacks across different models and datasets. Best results are shown in bold.
Model Defence LS LIE MM MS
MNIST (MLP) FedAvg 97.36 52.45 95.81 94.47
Mkrum 11.46 10.55 92.64 71.07
TrimMean 92.86 15.49 93.07 82.15
AFA 96.73 11.35 94.33 92.70
FLDetector 96.92 45.60 93.24 92.36
DeFL 97.47 88.90 94.29 93.29
CAD (Ours) 97.83 97.08 97.87 97.73
Fashion MNIST (AlexNet) FedAvg 88.06 15.66 78.71 74.58
Mkrum 10.02 10.00 27.13 18.61
TrimMean 75.17 10.00 56.02 53.31
AFA 88.56 10.00 23.70 23.70
FLDetector 76.51 10.00 78.37 74.98
DeFL 88.25 10.11 87.73 86.83
CAD (Ours) 88.79 87.86 88.43 88.32
Cifar10 (VGG-11) FedAvg 62.91 12.94 19.02 21.40
Mkrum 10.40 10.15 11.00 17.19
TrimMean 37.26 11.55 16.92 17.49
AFA 60.32 10.42 16.65 15.42
FLDetector 34.84 14.41 23.61 20.47
DeFL 63.02 23.17 21.60 48.26
CAD (Ours) 64.27 64.44 64.00 64.75
Cifar10 (AlexNet) FedAvg 59.63 10.16 19.86 19.75
Mkrum 11.77 10.00 17.61 16.83
TrimMean 33.80 10.00 19.34 17.44
AFA 57.02 10.01 18.68 19.47
FLDetector 32.47 10.00 17.24 17.17
DeFL 59.42 18.98 30.69 49.05
CAD (Ours) 62.76 62.90 62.92 62.60
@@ Conclusion
In this paper, we proposed a Confidence-Aware Defense (CAD) model, a novel defense mechanism for federated learning that achieve malicious detection to address both model poisoning and data poisoning attacks. The proposed CAD model estimates per-client confidence scores to reflect prediction uncertainty, which are then referred to identify clients into honest and malicious categories. By re-weighting the contributions of honest clients based on their confidence scores, CAD model enhances the robustness and accuracy of the global model. More importantly, CAD can effectively detect and mitigate various attack conditions by focusing on client confidence scores. Comprehensive experiments on multiple datasets and models demonstrate that CAD significantly improves the accuracy, stability, and robustness of federated learning systems compared to existing state-of-the-art methods.

### v2

@@ Introduction
Moderate (25%) FedAvg 55.40 64.37 49.44 59.20 49.84 52.89 68.32 68.60
TrimMean 53.69 61.35 51.58 60.33 31.63 34.22 65.20 65.20
AFA 68.89 68.71 68.99 69.17 69.09 68.80 68.88 69.09
FLDetector 32.66 32.37 25.93 35.61 61.94 56.92 39.74 46.32
DeFL 65.67 69.04 62.97 68.74 48.02 51.51 67.95 68.57
FedRoLA 69.23 69.03 68.59 68.76 68.67 68.95 68.48 68.59
CAD (Ours) 69.85 69.43 69.94 69.43 69.67 69.18 69.36 69.47
Severe (50%) FedAvg 19.02 18.37 21.40 22.25 12.94 15.04 62.91 62.70
TrimMean 16.92 17.27 17.49 17.20 11.55 12.35 37.26 36.88
AFA 16.65 17.89 15.42 17.42 10.42 10.56 60.32 60.85
FLDetector 23.61 20.02 20.47 20.73 14.41 11.96 34.84 38.81
DeFL 21.60 27.50 48.26 48.26 23.17 23.95 63.02 62.31
FedRoLA 19.52 19.60 21.40 22.10 13.42 14.17 62.52 61.98
CAD (Ours) 64.00 65.07 64.75 65.13 64.44 64.96 64.27 64.19
Extreme (75%) FedAvg 16.71 17.72 17.79 17.60 10.56 12.08 50.51 53.46
TrimMean 14.57 15.03 17.52 16.60 10.01 10.42 21.17 20.23
AFA 11.54 11.62 13.84 10.03 10.06 10.51 20.23 22.17
FLDetector 16.71 17.72 17.79 16.31 11.04 12.08 18.33 18.04
DeFL 20.11 19.44 19.42 20.02 16.06 16.76 52.18 52.72
FedRoLA 16.71 16.86 17.79 17.59 10.00 10.00 54.39 52.46
CAD (Ours) 56.96 57.04 56.88 56.48 56.75 56.54 57.68 57.79
Datasets.
We evaluated CAD model a wide range of benchmarks including: CIFAR-10 (Krizhevsky and Hinton 2009), MNIST (LeCun et al. 1998), and Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017) datasets, which are common in prior studies (Yan et al. 2023; Gang Yan and Li 2024). We employed a heterogeneous partitioning strategy to create non-IID local data where samples and class distributions are unbalanced among clients. The factor is sampled from a Dirichlet distribution $\texttt{Dir}_{N}(\alpha)$, with $\alpha$ determining the level of heterogeneity, where a smaller $\alpha$ value refers to high heterogeneity (Shang et al. 2022). We set $\alpha=0.5$ by default in the following experiments. We evaluated the performance by classification accuracy.
DNN Models.
We tested CAD model on three representative DNN models including: AlexNet (Krizhevsky, Sutskever, and Hinton 2012), VGG-11 (Simonyan and Zisserman 2015), and a Multi-Layer Perceptron (MLP) (Popescu et al. 2009) with layers consisting of 784, 512, and 10 neurons respectively. Specifically, AlexNet and VGG-11 were employed for the CIFAR-10 dataset, the MLP for MNIST, and AlexNet for Fashion-MNIST.
Baseline Defenses.
We compared CAD with state-of-the-art defense models including: TrimMean (Yin et al. 2018), AFA (Muñoz-González, Co, and Lupu 2019), FLDetector (Zhang and et al. 2022), DeFL (Yan et al. 2023), FedRoLA (Gang Yan and Li 2024). Additionally, we included FedAvg (McMahan et al. 2017) as a baseline method for comparison. We consider two settings regarding the adversary’s knowledge (Yan et al. 2023): (a) Full, where the adversary knows the gradients of honest clients, and (b) Partial, where the adversary is unaware of the gradient updates shared by honest clients. We evaluated the detection performance with True Positive Rate (TPR) and False Positive Rate (FPR).
Malicious Client Attacks.
We tested CAD model on both model poisoning attacks and data poisoning attacks:
@@
• Little is Enough (LIE): Adds small amounts of noise to the average of honest gradients. The adversary computes the average $\mu$ and standard deviation $\sigma$ of honest gradients, then applies a coefficient $z$ to obtain the malicious update $\mu+z\sigma$.
• Min-Max (MM): Computes malicious gradients such that their maximum distance from any gradient is upper bounded by the maximum distance between any two honest gradients, placing them close to the honest gradient cluster.
• Min-Sum (MS): The sum of squared distances of malicious gradients from all honest gradients is upper bounded by the sum of squared distances among honest gradients. Malicious gradients are kept identical to maximize impact.
Implementation Details.
We simulated $N=50$ clients in the experiments. Each client was trained for 20 epochs with stochastic gradient descent to update its local model. The number of training rounds was set to 200 to ensure convergence for all DNN models. The learning rate was set to 0.01. The batch size was set to 16, with a weight decay of $1\times 10^{-3}$. We evaluated the robustness of our CAD model under three degrees of attack severity defined by the percentage of malicious clients: moderate (25%), severe (50%), and extreme (75%). For data poisoning attacks, confidence scores were directly obtained during model training. For model poisoning attacks, confidence scores were obtained by further running inference on local training data before sending it to the parameter server. We implemented our defenses, in PyTorch (Paszke et al. 2019) on a single NVIDIA A100 GPU.
@@
Comparison on Defense Robustness under Different Attack Intensities.
@@
We evaluated the robustness of the proposed CAD model and other state-of-the-art defenses under different levels of attack intensity with the CIFAR-10 dataset with the VGG model. The results, as shown in Table 2, indicate that under moderate attacks (25% malicious clients), most defense methods perform reasonably well, with our CAD model achieving the highest accuracy across all attack types. However, as the intensity of the attacks increases to severe (50%) and extreme (75%), the performance of most methods significantly deteriorates. In contrast, the CAD model consistently maintains high accuracy and robustness by outperforming other methods under all attack intensities. This demonstrates the effectiveness of our confidence-based detection and re-weighted aggregation approach in mitigating the impact of various attack types, even under severe attack conditions.
Applicable to Various Models and Datasets.
@@
Table 3 compares the performance of CAD with other state-of-the-art defense methods across different models and datasets. The results demonstrate that CAD consistently outperforms other defenses with high accuracy under various attacks. For MNIST with MLP, most defense methods perform well under label shuffling attacks, but CAD achieves the highest accuracy across all attack types. Similarly, for Fashion MNIST with AlexNet, CAD maintains superior performance, especially against LIE and MM attacks, where other methods significantly collapsed in accuracy. In the CIFAR-10 dataset, both with VGG-11 and AlexNet models, the robustness of CAD stands out. While other defenses struggle to converge, particularly under LIE and MM attacks, CAD consistently achieves the best results. This highlights the effectiveness of our method in providing robust defense against various attacks, even under different model and dataset combinations. These results demonstrated that CAD is not only effective in mitigating attacks for a specific benchmark but also adaptable to various models and datasets, thereby making it a reliable defense method.
@@
Table 3: Defense robustness across various models and datasets. Accuracy of different defense methods with various attacks across different models and datasets under the severe attack setting. The best results are shown in bold.
Model Defence MM MS LIE LS
Cifar10 (VGG-11) FedAvg 19.02 21.40 12.94 62.91
AFA 16.65 15.42 10.42 60.32
FLDetector 23.61 20.47 14.41 34.84
DeFL 21.60 48.26 23.17 63.02
FedRoLA 19.52 21.40 13.42 62.52
CAD (Ours) 64.00 64.75 64.44 64.27
Cifar10 (AlexNet) FedAvg 19.86 19.75 10.16 59.63
AFA 18.68 19.47 10.01 57.02
FLDetector 17.24 17.17 10.00 32.47
DeFL 30.69 49.05 18.98 59.42
FedRoLA 19.52 19.75 10.16 59.33
CAD (Ours) 62.92 62.60 62.90 62.76
Fashion MNIST (AlexNet) FedAvg 78.71 74.58 15.66 88.06
AFA 23.70 23.70 10.00 88.56
FLDetector 78.37 74.98 10.00 76.51
DeFL 87.73 86.83 10.11 88.25
FedRoLA 78.09 74.58 15.66 88.05
CAD (Ours) 88.43 88.32 87.86 88.79
MNIST (MLP) FedAvg 95.81 94.47 52.45 97.36
AFA 94.33 92.70 11.35 96.73
FLDetector 93.24 92.36 45.60 96.92
DeFL 94.29 93.29 88.90 97.47
FedRoLA 95.81 94.47 52.45 97.40
CAD (Ours) 97.87 97.73 97.08 97.83
Robustness to Data Heterogeneity.
Figure 3: Robustness to data heterogeneity by varying degrees of Non-IID data, as controlled by the parameter $\alpha$ in the Dirichlet distribution $\texttt{Dir}(\alpha)$.
We further investigated the robustness of CAD across different levels of data heterogeneity by varying the key parameter $\alpha$ in the Dirichlet distribution $\text{Dir}(\alpha)$. A smaller $\alpha$ indicates a higher degree of non-IID data. We simulated a moderate LIE attack and a severe Label Shuffle attack on the Cifar10 dataset with VGG-11 model. As shown in Fig. 3, as the non-IID degree decreases (i.e., , as $\alpha$ increases), the global model accuracy improves at a much higher rate compared to FedAvg, as indicated by the blue shaded area, consistently under both Full and Partial scenarios. These results confirm the robustness of our CAD model to varying data heterogeneity.
Malicious Clients Detection Accuracy.
We evaluated the detection performance of our proposed CAD model under LIE and label shuffle attack scenarios considering both moderate and severe intensities. The evaluation was conducted on the Cifar10 dataset with the VGG-11 model at full attack setting. As shown in Fig. 4, our CAD model can accurately detect malicious clients, and converge to high TPR and low FPR after only a few training rounds. Therefore, it can exclude malicious clients and lead to a significant improvement in overall performance, as shown in the blue line over the orange line. Its advantage is more obvious when more clients are being attacked, with detection accuracy remaining high and stable even under severe attack conditions while the FedAvg was collapsed. This demonstrates that CAD is effective under varying intensities and thereby significantly enhances model performance compared to the FedAvg baseline.
Figure 4: Malicious client detection accuracy evaluated on the Cifar10 dataset with the VGG-11 model.
Impact of Re-weighting.
We ablated the impact of the re-weighting operation in Eq. (12) and the results are shown in Fig. 5. It can be observed that the re-weighting operation can consistently enhance performance, especially at higher attack rates. The improvement is particularly clear in “Severe” and “Extreme” conditions, with detecting attacks like LIE and MinSum benefiting the most. This effectiveness is likely due to the fact that only a few honest clients remain under high attack rates, therefore, it is crucial to identify and prioritize them. This further demonstrates that the confidence score is a reliable metric for assessing the quality of client models, thereby supporting the underlying assumption of our confidence-aware CAD model.
Figure 5: Effect of re-weighting on model accuracy across different degrees of attack.
@@ Conclusion
In this paper, we proposed the CAD model, an attack defense mechanism for FL that achieves accurate malicious client detection to address both model poisoning and data poisoning attacks. The proposed CAD model estimates per-client confidence scores to reflect the model’s uncertainty, which are then referred to identify clients into honest and malicious categories. By re-weighting the contributions of honest clients based on their confidence scores, CAD model can make the best use of honest clients even under severe attack when there is only a handful of honest clients available, so that it can enhance the robustness of the global model. Extensive experiments conducted on various benchmarks with non-IID data demonstrated that CAD can notably enhance the accuracy, stability, and robustness of FL systems.